# Implementation of the Martini-Ericson-Chanfray-Marteau RPA-based (anti)neutrino cross-section model in the GENIE neutrino event generator

### v1

@@ Abstract
We discuss the first implementation of the Martini-Ericson-Chanfray-Marteau RPA-based (anti)neutrino cross-section model for quasielastic (1p1h) and multinucleon (2p2h and 3p3h) excitations in the widely used GENIE neutrino event generator. Validation steps are presented, in particular, through direct comparisons of GENIE cross section output with original calculations performed by the authors of the model. Predictions for 12C, 16O and 40Ar are compared with some available T2K and MicroBooNE experimental measurements showing a reasonable agreement.
@@ I Introduction
The success of the next generation long-baseline accelerator-based neutrino oscillation experiments Hyper-Kamiokande Abe et al. (2018a) and DUNE Acciarri et al. (2015) will depend on the reduction to the percent level of the systematic errors. Presently, the most important source of uncertainties is represented by the neutrino-nucleus cross sections Katori and Martini (2018); Alvarez-Ruso et al. (2018), due to the difficulty in modeling nuclear dynamics, the nucleus being a quantum many-body system whose constituents interact through the strong force. To improve the knowledge of neutrino cross sections, a common and joint effort of experimental measurements, theoretical calculations and model implementations in Monte Carlo event generators is needed. These three aspects cannot be separated from each other and are entangled as three Borromean rings. Think, for example, of the neutrino energy reconstruction problem, the background subtraction for a peculiar experimental signal topology, or the modeling of the detector response.
@@
In this paper, we present the implementation of the theoretical RPA-based model of the Lyon group, described in detail in the article of Martini, Ericson, Chanfray and Marteau Martini et al. (2009) into the GENIE event generator Andreopoulos et al. (2010). We then compare its predictions with different experimental measurements. This model allows a unified description of several scattering channels: quasielastic, multinucleon emission and the coherent and incoherent single pion production. Within this model, the debated MiniBooNE quasielastic-like cross section on carbon Aguilar-Arevalo et al. (2010) has been explained for the first time Martini et al. (2009, 2010) by stressing the crucial role of the multinucleon (also called 2p2h or npnh) excitations contribution which sizably increase the cross section. After this explanation, 2p2h excitations started to attract a lot of attention (for a review see, for example, Refs. Katori and Martini (2018); Alvarez-Ruso et al. (2018)), first in the neutrino scattering community, then in the neutrino oscillation one. This growing interest is due to the central role of npnh, together with quasielastic processes, in the accurate reconstruction of the neutrino energy using the kinematic method employed by the T2K Abe et al. (2011) and Hyper-K experiments Martini et al. (2012, 2013); Nieves et al. (2012); Lalakulich et al. (2012); Ankowski et al. (2015), and consequently in the precise determination of the neutrino oscillation parameters. All widely-used event generators now include the multinucleon excitation channel, previously ignored before the MiniBooNE quasielastic-like results, into their packages. For a relatively recent state-of-the-art implementation in the different generators, see for example Refs. Alvarez-Ruso et al. (2021); Hayato and Pickering (2021); Golan et al. (2012); Buss et al. (2012), as well as the dedicated articles Schwehr et al. (2016); Dolan et al. (2020); Bourguille et al. (2021); Prasad et al. (2024). This implementation effort is still ongoing, essentially for two reasons. Firstly, various theoretical models present some differences in their predictions for this channel, whose complexity – stemming from the intrinsic difficulty of nuclear physics – makes it hard to treat without approximations. Secondly, the experimental neutrino community is moving towards more exclusive measurements based on detected hadronic variables, whereas most of the existing models were originally developed for inclusive cross sections in terms of leptonic observables.
@@
Up until now, the model of Ref. Martini et al. (2009) has not been included in any generators by the implementation of the corresponding theoretical hadron tensor, but only employing tables of shared cross sections at different kinematics via a re-weighting procedure Nguyen (2022). The hadron tensor strategy is employed here for the first time for this model. For a detailed explanation of this strategy, see Ref. Dolan et al. (2020), where it is introduced in the context of the SuSAv2 model. Even though, in principle, all the channels described by this model can be implemented in the generators, in the present work we confine ourselves to considering only the quasielastic (also called 1p1h) and the multinucleon excitations channel (2p2h and 3p3h, collectively referred to as npnh), which have been the most investigated and employed in the neutrino analyses of accelerator-based experiments over the last fifteen years. For this purpose we follow the same protocol as the one adopted in Refs. Dolan et al. (2020) and Dolan et al. (2022) for the SuSAv2+MEC Megias et al. (2016) (quasielastic + 2p2h) and Ghent CRPA Pandey et al. (2015) (giant resonances + quasielastic) models, respectively.
@@
The interest in implementing the Martini, Ericson, Chanfray, and Marteau model (hereafter referred to as the Martini et al. model) relies on several aspects. First, as already mentioned, differences appear among the various models, and none is currently able to simultaneously describe all the available experimental measurements across the full neutrino energy range – from low (e.g., T2K) to high (e.g., MINERvA) energies – and across nuclear targets (most notably, 12C, 16O and 40Ar). In the absence of such a solution, the span of the predictions from plausible models already implemented inside generators provides some guidance about, for example, the size of systematic errors associated to different processes (e.g., see the heavy use of this approach in Ref. Wilkinson et al. (2022)). Second, in the Martini et al. model the different 2p2h contributions – namely the ones related to i) nucleon-nucleon correlations, ii) meson exchange currents (MEC), and iii) their interference – as well as the 3p3h excitations, are separately calculated and available as distinct components of the hadron tensor. This allows for easier comparison between 2p2h models and, overall, could help avoid the risk of double counting (e.g., nucleon-nucleon correlations contributions) when combining different channels taken from different models (e.g., using a spectral function model, like that in Ref. Benhar and Meloni (2007), for the quasielastic channel and Fermi gas-based models for 2p2h interactions). Finally, calculations for different atomic nuclei are available, enabling the modeling of cross sections for carbon, water and argon targets – the main targets of present and future accelerator-based neutrino experiments.
@@ II The theoretical model
The theoretical model is described in Ref. Martini et al. (2009), to which we refer for all the details. It is based on nuclear response functions treated in the random phase approximation (RPA) on top of a local Fermi gas (LFG) calculation. This formalism belongs to a long tradition of investigations: first on electron scattering and pionic modes in nuclei (see, for example, Refs. Delorme et al. (1980); Alberico et al. (1980, 1982, 1984); Stroth et al. (1985); Alberico et al. (1985, 1986); Chanfray and Schuck (1988); Delorme and Guichon (1991); Laktineh et al. (1993); Chanfray et al. (1993); Alberico et al. (1998)), and later on neutrino scattering Delorme and Ericson (1985); Marteau (1999); Marteau et al. (2000). The model of Ref. Martini et al. (2009) is the continuation and extension of these works. One of its strengths is that it allows a unified description of several scattering channels: quasielastic, multinucleon emission and the coherent and incoherent single pion production. As already mentioned, within this model, the debated MiniBooNE quasielastic-like cross section on carbon Aguilar-Arevalo et al. (2010) was explained for the first time Martini et al. (2009, 2010) by advocating the crucial role of the multinucleon excitations. In Ref. Martini et al. (2011), an improvement of the original description of the quasielastic channel was made through the introduction of relativistic corrections. This relativistic description of the quasielastic channel is implemented in the present work. The Martini et al. model has successfully reproduced many of the features of several muon-neutrino flux-integrated double differential cross-section measurements in the past: the MiniBooNE Charged Current Quasielastic (CCQE) - like neutrino Martini et al. (2011) and antineutrino Martini and Ericson (2013) measurements, as well as their combination Ericson and Martini (2015); the MiniBooNE CC1 and the T2K CC inclusive measurements Martini and Ericson (2014); Martini et al. (2022); the T2K CC0 for neutrino Abe et al. (2016) and antineutrino measurements, as well as their combination Abe et al. (2020a). Single differential electron-neutrino T2K CC inclusive cross sections have been successfully reproduced as well Martini et al. (2016). More recently, the model has been employed to investigate the MicroBooNE CC inclusive Martini et al. (2022) and NC1 Martini et al. (2024a) cross sections on argon. The counter-intuitive dominance of the differential cross sections over the ones in specific kinematical conditions – not only for quasielastic, but also for multinucleon excitations and single pion production – has been explained with simple arguments in Ref. Martini et al. (2024b) using this model.
@@
Even though it has not yet been included in any Monte Carlo event generator via the implementation of the corresponding hadron tensor, the model has already been considered in many experimental neutrino cross-section measurement comparisons Abe et al. (2016, 2020a); Acero et al. (2020); Abe et al. (2023a) and oscillation analyses Abe et al. (2017, 2018b, 2020b, 2023b, 2025). In these experimental papers, the model is often referred to as “Martini et al.” or “Lyon”. The main reason for considering it is its larger cross-section predictions in the 2p2h sector (especially for neutrinos, and to a lesser extent for antineutrinos) compared to those of Ref. Nieves et al. (2011), a similar theoretical approach, often referred to as “Nieves et al.” or “Valencia”, which is already implemented in many generators. A very recent reexamination Sobczyk and Nieves (2025) of the Nieves et al. approach now leads to new results that are closer to that of Martini et al.
@@
Since the present work focuses mainly on multinucleon excitations, we recall that in the model of Ref. Martini et al. (2009) the following contributions are included: the nucleon-nucleon correlations (denoted NN in Ref. Martini et al. (2009)); the -MEC contribution (denoted in Ref.Martini et al. (2009)); and the NN correlations–-MEC interference, (denoted N in Ref. Martini et al. (2009)). We also recall that, as in all the works of the Lyon group starting from the end of Ref. Martini et al. (2009), only the npnh evaluation called “new” in Ref. Martini et al. (2009) is considered; the “old” version corresponds to the one employed in Refs. Marteau (1999); Marteau et al. (2000). This so-called “new” evaluation is based, for the NN and N parts, on the microscopic calculations of Ref. Alberico et al. (1984), aimed at determining the 2p2h contribution to the isospin spin-transverse response. For the part, the contributions related to the non-pionic decay of the inside the nuclear medium – leading to 2p2h and 3p3h excitations – are taken from Ref. Oset and Salcedo (1987), while the terms not reducible to a modifications of the width are taken from Ref. Delorme and Guichon (1991), which exploits the calculations of Ref. Shimizu and Faessler (1980). For further details, see Sec. IV C of Ref. Martini et al. (2009) and Sec. II of Ref. Martini et al. (2010). We also recall that all the form factors entering the hadron tensor coincide with those of Ref. Martini et al. (2009), with exactly the same parameters. For example, for the axial form factor, a standard dipole parameterization is adopted, with and .
@@ III Implementation in GENIE and validation
The model described above has been implemented in the GENIE neutrino event generator using the same approach as that adopted in Refs. Dolan et al. (2020) and Dolan et al. (2022). This approach relies on the fact that the differential cross section can be expressed as the product of kinematic factors and the contraction of a generic lepton tensor with a model-specific hadron tensor. The latter encodes all of the nuclear dynamics. In this framework, the implementation of the Martini et al. model is achieved by inserting new hadron tensor look-up tables into GENIE, as was previously done for other models in Refs. Schwehr et al. (2016), Dolan et al. (2020) and Dolan et al. (2022). These tables are provided for the 1p1h and npnh channels for carbon, oxygen and calcium targets. The npnh contribution includes both 2p2h and 3p3h components, and separate hadron tensors are employed for each. Despite the naming, the 3p3h implementation does not produce three final-state nucleons; instead, it is treated as an additional 2p2h-like contribution and results in two nucleons in the final state. Limitations of this approach are discussed in Sec. III.2. The input hadron tensors are defined over an energy transfer varying in the range , and a momentum transfer varying in the range , as in Ref. Martini et al. (2009). They are binned ( bins in energy transfer and bins in momentum transfer) and are evaluated using the same interpolation method of Refs. Dolan et al. (2020) and Dolan et al. (2022), similar to the one described in Ref. Schwehr et al. (2016).

### v2

# Implementation of the Martini-Ericson-Chanfray-Marteau RPA-based neutrino and antineutrino cross-section model in the GENIE neutrino event generator

@@ Abstract
We discuss the first implementation of the Martini-Ericson-Chanfray-Marteau random phase approximation-based (anti)neutrino cross-section model for quasielastic (1p1h) and multinucleon (2p2h and 3p3h) excitations in the widely used GENIE neutrino event generator. Validation steps are presented, in particular, through direct comparisons of GENIE cross-section output with original calculations performed by the authors of the model. Predictions for 12C, 16O, and 40Ar are compared with some available T2K and MicroBooNE experimental measurements showing a reasonable agreement.
@@ I Introduction
The success of the next generation long-baseline accelerator-based neutrino oscillation experiments Hyper-Kamiokande Abe et al. (2018a) and DUNE Acciarri et al. (2015a) will depend on the reduction to the percent level of the systematic errors. Presently, the most important source of uncertainties is represented by the neutrino-nucleus cross sections Katori and Martini (2018); Alvarez-Ruso et al. (2018), due to the difficulty in modeling nuclear dynamics, the nucleus being a quantum many-body system whose constituents interact through the strong force. To improve the knowledge of neutrino cross sections, a common and joint effort of experimental measurements, theoretical calculations and model implementations in Monte Carlo event generators is needed. These three aspects cannot be separated from each other and are entangled as three Borromean rings. Think, for example, of the neutrino energy reconstruction problem, the background subtraction for a peculiar experimental signal topology, or the modeling of the detector response.
@@
In this paper, we present the implementation of the theoretical random phase approximation (RPA) based model of the Lyon group, described in detail in the article of Martini et al. Martini et al. (2009) into the GENIE event generator Andreopoulos et al. (2010). We then compare its predictions with different experimental measurements. This model allows a unified description of several scattering channels: quasielastic (QE), multinucleon emission and the coherent and incoherent single pion production. Within this model, the debated MiniBooNE quasielasticlike cross section on carbon Aguilar-Arevalo et al. (2010) has been explained for the first time Martini et al. (2009, 2010) by stressing the crucial role of the multinucleon (also called 2p2h or npnh) excitation contributions which sizably increase the cross section. After this explanation, 2p2h excitations started to attract a lot of attention (for a review see, for example, Refs. Katori and Martini (2018); Alvarez-Ruso et al. (2018)), first in the neutrino scattering community, then in the neutrino oscillation one. This growing interest is due to the central role of npnh, together with quasielastic processes, in the accurate reconstruction of the neutrino energy using the kinematic method employed by the T2K Abe et al. (2011) and Hyper-K experiments Martini et al. (2012, 2013); Nieves et al. (2012); Lalakulich et al. (2012); Ankowski et al. (2015) and, consequently, in the precise determination of the neutrino oscillation parameters. All widely-used event generators now include the multinucleon excitation channel, previously ignored before the MiniBooNE quasielasticlike results, into their packages. For a relatively recent state-of-the-art implementation in the different generators see, for example, Refs. Alvarez-Ruso et al. (2021); Hayato and Pickering (2021); Golan et al. (2012); Buss et al. (2012), as well as the dedicated articles Schwehr et al. (2016); Dolan et al. (2020); Bourguille et al. (2021); Prasad et al. (2024); Mosel and Gallmeister (2024); Ankowski et al. (2025). This implementation effort is still ongoing, essentially for two reasons. First, various theoretical models present some differences in their predictions for this channel, whose complexity–stemming from the intrinsic difficulty of nuclear physics–makes it hard to treat without approximations. Second, the experimental neutrino community is moving toward more exclusive measurements based on detected hadronic variables, whereas most of the existing models were originally developed for inclusive cross sections in terms of leptonic observables. This interest in the hadronic variables is driven by the improved sensitivity to outgoing low-energy protons and neutrons of the upgraded T2K near detector ND280 Abe et al. (2019) as well as by the present Short-Baseline Neutrino Program at Fermilab Acciarri et al. (2015b) and the future DUNE experiment Acciarri et al. (2015a), both of which use liquid argon detectors with especially good hadronic sensitivity. The calorimetric energy estimator used there can also be problematically sensitive to the npnh effects (e.g. missing energy due to emitted neutrons or subthreshold protons).
@@
Up until now, the model of Ref. Martini et al. (2009) has not been included in any generators by the implementation of the corresponding theoretical hadron tensor, but only employing tables of shared cross sections at different kinematics via a reweighting procedure Nguyen (2022). The hadron tensor strategy is employed here for the first time for this model. For a detailed explanation of this strategy, see Ref. Schwehr et al. (2016) where it was developed as well as Ref. Dolan et al. (2020), where it is discussed in the context of the SuSAv2 model. Even though, in principle, all the channels described by this model can be implemented in the generators, in the present work we confine ourselves to considering only the quasielastic (also called 1p1h) and the multinucleon excitation channels (2p2h and 3p3h, collectively referred to as npnh), which have been the most investigated and employed in the neutrino analyses of accelerator-based experiments over the last fifteen years. For this purpose we follow the same protocol as the one adopted in Refs. Dolan et al. (2020, 2022) for the SuSAv2+ meson-exchange current (MEC) Megias et al. (2016) (quasielastic + 2p2h) and Ghent continuum random-phase approximation (CRPA) Pandey et al. (2015) (giant resonances + quasielastic) models, respectively.
@@
The interest in implementing the Martini-Ericson-Chanfray-and Marteau model (hereafter referred to as the Martini et al. model) relies on several aspects. First, as already mentioned, differences appear among the various models, and none is currently able to simultaneously describe all the available experimental measurements across the full neutrino energy range–from low (e.g., T2K) to high (e.g., MINERvA) energies–and across nuclear targets (most notably, 12C, 16O, and 40Ar). In the absence of such a solution, the span of the predictions from plausible models already implemented inside generators provides some guidance about, for example, the size of systematic errors associated with different processes (e.g., see the heavy use of this approach in Ref. Wilkinson et al. (2022)). Second, in the Martini et al. model, the different 2p2h contributions–namely the ones related to (i) nucleon-nucleon correlations, (ii) meson exchange currents, and (iii) their interference–as well as the 3p3h excitations, are separately calculated and available as distinct components of the hadron tensor. This allows for easier comparison between 2p2h models and, overall, could help avoid the risk of double counting (e.g., nucleon-nucleon correlations contributions) when combining different channels taken from different models (e.g., using a spectral function model, like that in Ref. Benhar and Meloni (2007), for the quasielastic channel and Fermi gas-based models for 2p2h interactions). Finally, calculations for 12C, 16O and 40Ca are available, enabling the modeling of cross sections for carbon, water, and argon targets–the main targets of present and future accelerator-based neutrino experiments.
Even if the Martini et al. model allows also the treatment of neutral-current (NC) neutrino cross sections, and the nuclear responses are available also for the electron scattering, in the present work we focus exclusively on the GENIE implementation of the neutrino and antineutrino (1p1h and npnh) charged-current (CC) cross sections.
@@ II The theoretical model
The theoretical model is described in Ref. Martini et al. (2009), to which we refer for all the details. It is based on nuclear response functions treated in the RPA on top of a local Fermi gas (LFG) calculation. This formalism belongs to a long tradition of investigations: first on electron scattering and pionic modes in nuclei (see, for example, Refs. Delorme et al. (1980); Alberico et al. (1980, 1982, 1984); Stroth et al. (1985); Alberico et al. (1985, 1986); Chanfray and Schuck (1988); Delorme and Guichon (1991); Laktineh et al. (1993); Chanfray et al. (1993); Alberico et al. (1998)) and, later on, neutrino scattering Delorme and Ericson (1985); Marteau (1999); Marteau et al. (2000). The model of Ref. Martini et al. (2009) is the continuation and extension of these works. One of its strengths is that it allows a unified description of several scattering channels: quasielastic, multinucleon emission and the coherent and incoherent single pion production. As already mentioned, within this model, the debated MiniBooNE quasielasticlike cross section on carbon Aguilar-Arevalo et al. (2010) was explained for the first time Martini et al. (2009, 2010) by advocating the crucial role of the multinucleon excitations. In Ref. Martini et al. (2011), an improvement of the original description of the quasielastic channel was made through the introduction of relativistic corrections. This relativistic description of the quasielastic channel is implemented in the present work. The Martini et al. model has successfully reproduced many of the features of several muon-neutrino flux-integrated double differential cross-section measurements in the past: the MiniBooNE charged-current quasielasticlike neutrino Martini et al. (2011) and antineutrino Martini and Ericson (2013) measurements, as well as their combination Ericson and Martini (2015); the MiniBooNE CC1 and the T2K CC inclusive measurements Martini and Ericson (2014); Martini et al. (2022); the T2K CC0 for neutrino Abe et al. (2016) and antineutrino measurements, as well as their combination Abe et al. (2020a). Single differential electron-neutrino T2K CC inclusive cross sections have been successfully reproduced as well Martini et al. (2016). More recently, the model has been employed to investigate the MicroBooNE CC inclusive Martini et al. (2022) and NC1 Martini et al. (2024a) cross sections on argon. The counterintuitive dominance of the differential cross sections over the ones in specific kinematical conditions–not only for quasielastic, but also for multinucleon excitations and single pion production–has been explained with simple arguments in Ref. Martini et al. (2024b) using this model.
@@
Even though it has not yet been included in any Monte Carlo event generator via the implementation of the corresponding hadron tensor, the model has already been considered in many experimental neutrino cross-section measurement comparisons Abe et al. (2016, 2020a); Acero et al. (2020); Abe et al. (2023a) and oscillation analyses Abe et al. (2017, 2018b, 2020b, 2023b, 2025). In these experimental papers, the model is often referred to as “Martini et al.” or “Lyon.” The main reason for considering it is its larger cross-section predictions in the 2p2h sector (especially for neutrinos and, to a lesser extent, for antineutrinos) compared to those of Ref. Nieves et al. (2011), a similar theoretical approach, often referred to as “Nieves et al.” or “Valencia,” which is already implemented in many generators. A very recent reexamination Sobczyk and Nieves (2025) of the Nieves et al. approach now leads to new results that are closer to that of Martini et al.
Before concluding this section, for clarity and to make the work as self-contained as possible (without entering in too much detail already published elsewhere) we remind the reader of some general aspects of the neutrino-nucleus cross sections and of the formalism employed by Martini et al. in their calculations.
The double differential cross section for the charged-current reaction is given by
Here is the differential solid angle in the direction specified by the charged lepton momentum in the laboratory frame, is the energy transferred to the nucleus, the zero component of the four-momentum transfer , with and , being the initial and final lepton four momenta. In Eq. (1), is the weak coupling constant, is the Cabibbo angle.
is the leptonic tensor
and is the hadronic tensor, a combination of nucleon form factors (that contain the information about the nucleon properties), and nuclear response functions (that contain the information about the nuclear dynamics). The different components of the hadronic tensor can be combined allowing a reformulation of Eq. (1) in terms of projections with respect to the momentum transfer direction. The charged-current cross section is a linear combination of five contributions
where the kinematical factors come from the contraction with the leptonic tensor and the plus (minus) sign applies to neutrinos (antineutrinos). We do not provide here the explicit expressions for the kinematical factors and for the nucleon form factors and nuclear responses entering in the hadronic tensors, as they can be found in Appendixes A and B of Ref. Martini et al. (2009).111There is a typo in the expression of in Eq. (A4) of Ref. Martini et al. (2009): the factor 8 should be replaced by the factor 16.
The hadronic tensor can be decomposed in terms of the following scattering channels:
that are available as subtables of tensor elements. It can also be decomposed in terms of the different channels of particle-hole (nucleon or Delta ) excitations
which are particularly useful to classify the type of 2p2h contributions. It is important to remember that Eqs. (1) and (3) are totally general and, according to the decompositions of Eq. (4), apply to different excitations channels: 1p1h quasielastic, 2p2h, 3p3h (the three channels considered in the present work), and 1p1h 1 production, coherent production (the two other channels of the Martini et al. model, whose Monte Carlo implementation is left for future works).
@@
Since the present work focuses mainly on multinucleon excitations, we recall that in the model of Ref. Martini et al. (2009) the following contributions are included: the nucleon-nucleon correlations (denoted NN in Ref. Martini et al. (2009)); the -MEC contribution (denoted in Ref Martini et al. (2009)); and the NN correlations–-MEC interference (denoted N in Ref. Martini et al. (2009)). These contributions are separately available as subtable of the hadron tensor , according to the decompositions of Eq. (5) , which holds for the different scattering channels. We also recall that, as in all the works of the Lyon group starting from the end of Ref. Martini et al. (2009), only the npnh evaluation called “new” in Ref. Martini et al. (2009) is considered; the “old” version corresponds to the one employed in Refs. Marteau (1999); Marteau et al. (2000). This so-called new evaluation is based, for the NN and N parts, on the microscopic calculations of Ref. Alberico et al. (1984), aimed at determining the 2p2h contribution to the isospin spin-transverse response. For the part, the contributions related to the nonpionic decay of the inside the nuclear medium–leading to 2p2h and 3p3h excitations–are taken from Ref. Oset and Salcedo (1987), while the terms not reducible to modifications of the width are taken from Ref. Delorme and Guichon (1991), which exploits the calculations of Ref. Shimizu and Faessler (1980). Further details are given in Appendix B. We also recall that, for the purpose of validating the present implementation against the previously published theoretical results, all the form factors entering the hadron tensor coincide with those of Ref. Martini et al. (2009), with exactly the same parameters. For example, for the axial form factor, a standard dipole parametrization is adopted, with and the value222We remind the reader that the PDG 2024 Navas et al. (2024) value of the axial coupling constant is slightly different, being . The effects of this difference with respect to our adopted value are expected to be small. They could be quantified in a future work. of .
@@ III Implementation in GENIE and validation
The model described above has been implemented in the GENIE neutrino event generator using the same approach as that adopted in Refs. Dolan et al. (2020, 2022). This approach relies on the fact that the differential cross section can be expressed as the product of kinematic factors and the contraction of a generic lepton tensor with a model-specific hadron tensor. The latter encodes all of the nuclear dynamics. In this framework, the implementation of the Martini et al. model is achieved by inserting new hadron tensor look-up tables into GENIE, as was previously done for other models in Refs. Schwehr et al. (2016); Dolan et al. (2020, 2022). These tables are provided for the 1p1h and npnh channels for carbon, oxygen, and calcium targets. The npnh contribution includes both 2p2h and 3p3h components, and separate hadron tensors are employed for each. Despite the naming, the 3p3h implementation does not produce three final-state nucleons; instead, it is treated as an additional 2p2h-like contribution and results in two nucleons in the final state. Limitations of this approach are discussed in Sec. III.2. The input hadron tensors are defined over an energy transfer varying in the range 5 and a momentum transfer varying in the range 1 , as in Ref. Martini et al. (2009). They are binned ( bins in energy transfer and bins in momentum transfer) and are evaluated using the same interpolation method of Refs. Dolan et al. (2020, 2022), similar to the one described in Ref. Schwehr et al. (2016).